OpenStudy (anonymous):

According to the natural sequence of exponents, how would you explain 2^-3?

OpenStudy (amistre64):

r^0 = 1 r^1 = r r^2 = rr r^3 = rrr this is a geometric sequence, \(r^n=r^{n-1}~r\) \[r^n=r^n~r^{-1}~r\] \[1=r^{-1}~r\] \[r^{-1}~r=1\] \[r^{-1}=\frac 1r\]

OpenStudy (amistre64):

therefore:\(2^{-3}=1/2^3\)

OpenStudy (amistre64):

something like that :)

OpenStudy (anonymous):

thank u!(:

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

thank u again :3